Ask
your own question, for FREE!
Mathematics
14 Online
A chef draws cookies randomly from a box containing 6 cookies of the same shape and size. There is 1 chocolate cookie, 3 almond cookies, and 2 butter cookies. He draws 1 cookie and then draws another cookie without replacing the first one. Find the probability of picking 1 almond cookie followed by another almond cookie, and show the equation used.
Still Need Help?
Join the QuestionCove community and study together with friends!
1 + 3 + 2 = 6 1 / 6 = 0.16 x 100 = 16% 1 / 5 = 0.2 x 122 = 20%
ok so he first has a chance of 3/6=1/2 for the almond cookie then after that he has 5 cookies left and so has a 2/5 chance to find the probabilty of one after another you have to multiply 3/6*2/5
@tonycoolkid21 wrote:
1 + 3 + 2 = 6
1 / 6 = 0.16 x 100 = 16%
1 / 5 = 0.2 x 122 = 20%
ok
ok
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
10 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal